# Ensemble Kalman Diffusion Guidance: A Derivative-free Method for Inverse Problems

### v1

@@ A.3 Zero-order gradient estimation baseline
Posterior sampling requires computation of the scores $\nabla_{{\bm{x}}_{t}}p({\bm{x}}_{t})$ and $\nabla_{{\bm{x}}_{t}}p({\bm{y}}\mid{\bm{x}}_{t})$; the former is learned by the pre-trained diffusion model, and the latter can be estimated by various approximation methods. In our baseline derivative-free inverse problem solver, we substitute the explicit automatic differentiation used in algorithms such as DPS with (28) and (29). We estimate this gradient by leveraging the fact that a probability flow ODE deterministically maps every ${\bm{x}}_{t}$ to ${\bm{x}}_{0}$; ^$\hat{\nabla}_{\hat{{\bm{x}}}_{0}}p({\bm{y}}\mid\hat{{\bm{x}}}_{0})$ is approximated with Gaussian smoothing, and a vector-Jacobian product (VJP) is used to then calculate ^$\hat{\nabla}_{{\bm{x}}_{t}}p({\bm{y}}\mid{\bm{x}}_{t})$. Our gradient estimate is defined as follows:
@@
${\bm{D}}_{{\bm{x}}_{t}}^{\top}$ is the transpose of the Jacobian matrix; (30) can be efficiently computed using automatic differentiation. Note that although automatic differentiation is used, differentiation through the forward model does not occur. Thus, this method is still applicable to non-differentiable inverse problems. Furthermore, we choose to perturb $\hat{{\bm{x}}}_{0}$ and use a VJP rather than directly perturb ${\bm{x}}_{t}$ so that we can avoid repeated forward passes through the pre-trained network, which is very expensive. Pseudocode for these algorithms is provided in Algorithm LABEL:alg:dps-zo.
@@ A.4 EnKG Implementation Details
There are mainly two design choices in our algorithm 2 to be made. The first is the step size $w_{i}$ which controls the extent to which the correction step moves towards the MAP estimator. In the ensemble Kalman literature (Kovachki & Stuart, 2019), the following adaptive step size is widely used and we adopt it for our experiments as well.
@@ A.6 Additional results
Figure 6: Ablation study on the number of particles for Navier-Stokes. The shaded region represents best and worst particle.
@@
Table 6: Hyperparameter choices for DPS with zero-order gradient estimation ($64\times 64$).
@@
Table 7: Hyperparameter choices for baselines based on zero-order gradient estimation ($256\times 256$).

### v2

@@ A.3 Zero-order gradient estimation baseline
Posterior sampling requires computation of the scores $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t})$ and $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}\mid{\bm{x}}_{t})$; the former is learned by the pre-trained diffusion model, and the latter can be estimated by various approximation methods. In our baseline derivative-free inverse problem solver, we substitute the explicit automatic differentiation used in algorithms such as DPS with (37) and (38). We estimate this gradient by leveraging the fact that a probability flow ODE deterministically maps every ${\bm{x}}_{t}$ to ${\bm{x}}_{0}$; ^$\hat{\nabla}_{\hat{{\bm{x}}}_{0}}\log p({\bm{y}}\mid\hat{{\bm{x}}}_{0})$ is approximated with Gaussian smoothing, and a vector-Jacobian product (VJP) is used to then calculate ^$\hat{\nabla}_{{\bm{x}}_{t}}\log p({\bm{y}}\mid{\bm{x}}_{t})$. Our gradient estimate is defined as follows:
@@
${\bm{D}}_{{\bm{x}}_{t}}^{\top}$ is the transpose of the Jacobian matrix; (39) can be efficiently computed using automatic differentiation. Note that although automatic differentiation is used, differentiation through the forward model does not occur. Thus, this method is still applicable to non-differentiable inverse problems. Furthermore, we choose to perturb $\hat{{\bm{x}}}_{0}$ and use a VJP rather than directly perturb ${\bm{x}}_{t}$ so that we can avoid repeated forward passes through the pre-trained network, which is very expensive. Pseudocode for these algorithms is provided in Algorithm 3.
@@ A.4 EnKG Implementation Details
There are mainly two design choices in our algorithm 2 to be made. The first is the step size $w_{i}$ which controls the extent to which the correction step moves towards the MAP estimator. In the ensemble Kalman literature (Kovachki & Stuart, 2019), the following adaptive step size is widely used, and we adopt it for our experiments as well.
@@ A.6 Additional results
Figure 6: Ablation studies on the impact of ensemble size, guidance scale, and the number of diffusion model queries, conducted on the Navier-Stokes problem. Left: Relative $L^{2}$ error versus ensemble size; the shaded region indicates the range between the best and worst particle. Middle: Relative $L^{2}$ error versus the guidance scale. Right: Relative $L^{2}$ error versus the number of diffusion model queries per particle; error bars indicate one standard deviation.
@@
Table 6: Hyperparameter choices for Forward-GSG and Central-GSG ($64\times 64$).
@@
Table 7: Hyperparameter choices for baselines Forward-GSG and Central-GSG ($256\times 256$).
@@ A.7 Details of black hole imaging
A.8 Additional comparison
To provide a more comprehensive evaluation, we provide comparisons against several gradient-based methods across different tasks.
Image restoration on FFHQ256
Table 8 presents comparisons with DPS (Chung et al., 2023b) and DiffPIR (Zhu et al., 2023) on four image restoration tasks: inpainting, super-resolution (x4), deblurring, and phase retrieval. We observe that EnKG achieves performance comparable to gradient-based methods, with no single approach emerging as a clear winner across all tasks. This demonstrates that EnKG offers competitive performance while maintaining its derivative-free property.
Navier-Stokes equation
Table 9 reports comparison with DPS and PnP-DM (Wu et al., 2024) on the Navier-Stokes equation problem. We observe that EnKG clearly outperforms the gradient-based methods, while PnP-DM encounters numerical instability, resulting in either a crash or timeout. Since DPS and PnP-DM do not have such experiments in their paper, we perform a grid search for its guidance scale over range $[10^{-3},10^{2}]$ to find the best choice. For PnP-DM, we explore all hyperparameter combinations mentioned in their paper; however, all result in a numerical crash within the PDE solver. Although reducing the Langevin Monte Carlo learning rate improved stability, it led to infeasible runtimes (e.g., exceeding 100 hours). Consequently, we mark PnP-DM as “crashed/timeout” in Table 9. Additionally, in this problem, autograd encounters out-of-memory issues when the pseudospectral solver unrolls beyond approximately 6k steps on an A100-40GB GPU. This limitation suggests that gradient-based methods may not be applicable to more complex problems that require a large number of PDE solver iterations.
Black hole imaging
As shown in Table 10 shows additional comparisons for the black hole imaging problem, including DPS and PnP-DM. Once again, EnKG delivers performance comparable to gradient-based methods. For DPS, we performed a grid search to optimize hyperparameters, while for PnP-DM, we used the settings provided in their paper. These results further demonstrate the robustness and competitiveness of EnKG across diverse scientific inverse problems.
Discussion
As shown in the experiments above, gradient-based methods do not always perform better than derivative-free methods. We believe this is due to two reasons. First, gradient-based methods are more prone to getting stuck in local optima, especially in noisy, non-smooth settings. In contrast, many derivative-free methods, including EnKG, incorporate implicit smoothing, which can improve robustness in such cases. Second, gradient-based methods often rely on Gaussian approximations via Tweedie’s formula, as taking gradients through the full unrolled reverse process is computationally infeasible. In contrast, EnKG gets a better likelihood approximation by simulating the reverse process without the need of backpropagation. More recent gradient-based works, such as Rout et al. (2024a); Chung et al. (2024), focus on solving linear inverse problems with text-to-image latent diffusion models. In contrast, EnKG does not rely on text-to-image latent diffusion model, which is generally unavailable in many inverse problem applications.
Table 8: Additional comparison with a few gradient-based methods on FFHQ 256x256 dataset. We report average metrics for image quality and consistency on four tasks. Measurement noise is $\sigma=0.05$ unless otherwise stated.
Inpaint (box) SR ($\times$4) Deblur (Gauss) Phase retrieval
[…]
Gradient-based
DiffPIR 22.87 0.653 0.268 26.48 0.744 0.220 24.87 0.690 0.251 22.20 0.733 0.270
Black-box access
Table 9: Additional comparison of relative L2 error on the Navier-Stokes inverse problem. Numbers in parentheses represent the sample standard deviation.
$\sigma_{\mathrm{noise}}=0$ $\sigma_{\mathrm{noise}}=1.0$ $\sigma_{\mathrm{noise}}=2.0$
Gradient-based
DPS 0.308 (0.214) 0.349 (0.246) 0.382 (0.228)
PnP-DM Crashed or timeout Crashed or timeout Crashed or timeout
Black-box access
EnKG(Ours) 0.120 (0.085) 0.191 (0.057) 0.294 (0.061)
Table 10: Additional comparison with a few gradient-based methods on the black-hole imaging problem.
DPS 23.984 26.220 1.212 1.079
PnP-DM 28.211 32.499 1.120 1.224
EnKG (Ours) 29.093 32.803 1.426 1.270
A.9 Robustness to the pretrained prior quality
In this section, we conduct a controlled experiment on Navier-Stokes equation problem to investigate the performance dependence on the quality of pre-trained diffusion models. Specifically, we trained a diffusion model prior using only 1/10 of the original training set and limited the training to 15k steps to simulate a lower-quality model. We evaluate the top two algorithms, EnKG and DPG, with the same hyperparameters used in the main experiments.
Robust performance
As shown in Table 11, we observe that while both algorithms experienced a performance drop due to the reduced quality of the diffusion model, the decline was relatively small compared to the significant reduction in training data. Notably, our EnKG demonstrated greater robustness, with a smaller performance drop than the best baseline method, DPG. These results indicate that while EnKG benefits from high-quality diffusion models, it is not overly sensitive to their quality. It maintains strong performance even with reduced model capabilities.
Table 11: Relative L2 error of DPG and EnKG (ours) with different diffusion model quality.
Original model trained with full data New model trained with $1/10$ data
DPG 0.325 (0.188) 0.394 (0.178)
EnKG (Ours) 0.120 (0.085) 0.169 (0.117)
Table 12: Per-particle computational resource usage of EnKG across three inverse problems. We use 1024 particles for FFHQ and black hole imaging, and 2048 particles for Navier-Stokes. NFE-DM: number of diffusion model evaluations. NFE-Fwd: number of forward model evaluations. Runtime is averaged per particle.
NFE-DM NFE-Forward Runtime (s) Peak GPU memory (GB)
FFHQ 256$\times$256 1632 144 11.6 23.9
Navier-Stokes equation 1632 144 3.6 7.1
Black hole imaging 771 60 1.1 1.4